The price of 5 kg of sweets was rs.180. The price has gone up such that the price of 4 kgs of sweet is now rs.180. What is the percentage increase in price?
Answer by mananth(16949)

5kg ---Rs180
1kg=36
.....
4kg-----Rs.180
1 kg =45
increase = 45-36=9
9/36=x/100
100*9/36
=25% increase